# thanks: Now at the European Space Technology Centre, European Space Agency, 2200 AG Noordwijk, The Netherlands

### v1

@@ III Simulator and simulation settings
For the fit and subtraction of the TTL coupling, we will use the scheme first introduced in [7] and further improved in [8]. We consider simulated data for one day downsampled to 0.5 Hz. First, we compute the Fourier transformations of the TDI AET variables of the interferometric length measurements including the TTL coupling noise. The same is done for the TDI AET variables of the angular couplings into the length measurements, i.e. the 12 measurements $\alpha_{ij},\ \alpha\in\{\eta,\varphi\},\ i\neq j\in\{1,2,3\}$ and their 12 delays along the corresponding arm. For the derivation of the angular coupling terms, we effectively propagate equation (3) through TDI, setting all coupling coefficients but one to zero. An MCMC algorithm is then used to fit the coupling coefficients to these two data sets. Like in [7, 8], we chose the TTL coefficients to be 2.3 mm/rad in all test cases. All fits were applied in the frequency range from 3 mHz to 0.9 Hz. We whitened the data iteratively. Therefore, we considered the LISA displacement noise and TM acceleration noise, but no prior knowledge of the GW signals included in the simulated data. Further details on the fitting algorithm and the data preparation are given in [8].
In [7], the LISASim simulator [21] was used to generate the interferometer data. This simulator does not yet include GW signals that we want to study here. Therefore, we used the LISANode simulator [22, 23], which computes the response of given GW signals in dependence on the orbital position of the SC. In LISANode, the length differences are given as frequency fluctuations, unlike LISASim, where they are given as length variations. To use the same code as in [7], we converted the frequency data into length data. Even though LISANode simulates anti-alias filters, we applied an additional low-pass filter to enhance the robustness. The same jitter levels and shapes have been implemented in both simulators. Moreover, only minor differences exist for the instrument noise contributors, such that we find an overall comparable level of instrument noise in the data from both simulators. The jitter and noise shapes used in our work are summarized in appendix A. We also find no significant differences in the accuracy of the TTL coupling coefficient estimation and the level of residual noise after TTL coupling subtraction. A more detailed comparison of the results using either of the two simulators is given in [8].
@@ IV TTL coupling estimation for changing arm lengths
In figure 2 we show the deviations of the estimated coupling coefficients from the true value. We find small differences between the estimates in the different cases. All estimation errors are below 0.06 mm/rad and therefore below the requirement of 0.1 mm/rad. To demonstrate the performance of the TTL noise subtraction, we plot the residuals in figure 3. We find that the residuals after subtraction in the cases of the static (light-red) and the ESA science orbit (blue) are almost identical. We did not plot the residual for the Keplerian orbit case since it looks like the other two residual curves. This finding supports the assumption in [7] that it is sufficient to consider static unequal arm lengths in TTL coupling subtraction analyses. We also plot the TTL noise residual (green) for the scenario with ESA science orbits. In the present paper, the term TTL noise residual describes the non-suppressed TTL noise and the DWS noise added due to the subtraction [7, cf. eq. (18)]. It is well below the other instrument noises.
@@ V.1 Verification binaries
As an initial test of the TTL coupling coefficient estimation in the presence of GW signals, we include the response of two verification binaries. The signal characteristics of such binary systems are well studied and can be easily distinguished from the TTL coupling noise due to their distinct response at a single frequency. Therefore, we use them to verify our correct inclusion of the signals in the simulated data. We chose the verification binaries HMCnc and ZTFJ1539 for our analysis. Their orbital frequencies are 6.220 mHz (HMCnc) and 4.822 mHz (ZTFJ1539) [31]. While most of the verification binaries listed in [31] enter the measurement band below the frequency range where TTL coupling is expected to dominate, the selected binaries lie in the frequency regime considered in our fit.
Figure 4: Performance of the TTL coupling subtraction in the scenario with verification binaries. The ASD of the TDI X combination for the simulations with TTL coupling is plotted in black. The dark-blue curve shows the ASD of the noise residual after the subtraction of the TTL coupling using the MCMC fit results and the light-blue dashed curve for the subtraction using the true coupling coefficients. The residual TTL coupling is plotted in dash-dotted green. The ASD of the GW response in TDI X is shown in purple. All ASDs were computed for one day measurements at 0.5 Hz. The dashed gray line corresponds to the LISA requirement.
@@ V.3 Detached white dwarf galactic binaries
Figure 9: Close-up view on the performance of the TTL coupling subtraction in the case GB1. The ASD of the TDI X combination for the simulations with (without) TTL coupling is plotted in black (gray). The dashed dark-blue curve shows the ASD of the residual after the subtraction of the TTL coupling using the MCMC fit results. The ASD of the GW response after TDI is shown in purple. The ASD of the residual after subtracting the TTL model with the fitted MCMC coefficients from data without instrument noise is plotted in dashed yellow. All ASDs were computed for one day measurements at 0.5 Hz.
@@ VI Summary
In the present paper, we have investigated the performance of the TTL coupling suppression via subtraction in post-processing for different GW signals in the simulated data. The data were generated using the simulator LISANode, and we chose the same noise realizations in all scenarios for a better comparison of the different cases. We then used an MCMC algorithm to fit the TTL coupling coefficients to the TDI AET combinations of the interferometric length measurements. Our analysis showed that the coupling coefficients could be fitted to within a deviation of $\pm$0.1 mm/rad in all the scenarios. By subtracting the TTL noise model from the data using the fitted coupling coefficients, we could suppress the total noise below the LISA mission noise requirement. The residual TTL noise was shown to be well below the other instrument noises. Due to the small deviations of the coupling coefficients, the differences of the residuals after subtraction using the fitted or the true coupling coefficients are marginal. We also investigated the GW signals after the TTL coupling subtraction. We could show that neither the coefficient inaccuracies nor the DWS readout noise added significantly to the data when subtracting the TTL noise. These results are a first evidence that the TTL suppression scheme planned for the LISA mission will not only suppress the noise well below the requirement, but will also leave the GW signals unchanged for successful source investigations.
Despite the overall reassuring results of our analysis, some aspects will require further investigation in the future. Our simulations of the LISA data considered only the SC and MOSA jitter, which is white in the frequency range where TTL coupling is assumed to be dominant. In reality, we expect the jitter to be colored. For instance, the SC jitter will roll off at higher frequencies reducing the frequency range in which TTL coupling is dominant. We have tested the TTL coefficient fit for smaller frequency ranges than the one used in most of this paper and found comparable results (see e.g. appendix B). It remains to be shown that the GW signals and the jitter can also be well distinguished for other jitter shapes. The implementation of the LISA dynamics and control loops [38, 39] in the LISANode simulator will allow more realistic tests in the future. Furthermore, our best estimates of instrument noise are constantly changing and may require updates of some of the presented tests. Finally, we have not yet considered all possible GW source types. E.g., we have not yet investigated extreme mass ratio inspirals or the cosmological SGWB.

### v3

@@ III Simulator and simulation settings
For the fit and subtraction of the TTL coupling, we will use the scheme first introduced in [7] and further improved in [8]. We consider simulated data for one day. First, we compute the Fourier transformations of the TDI AET variables of the interferometric length measurements including the TTL coupling noise. The same is done for the TDI AET variables of the angular couplings into the length measurements, i.e. the 12 measurements $\alpha_{ij},\ \alpha\in\{\eta,\varphi\},\ i\neq j\in\{1,2,3\}$ and their 12 delays along the corresponding arm. For the derivation of the angular coupling terms, we effectively propagate equation (3) through TDI, setting all coupling coefficients but one to zero. An MCMC algorithm is then used to fit the coupling coefficients to these two data sets. Like in [7, 8], we chose the TTL coefficients to be 2.3 mm/rad in all test cases. All fits were applied in the frequency range from 3 mHz to 0.9 Hz. We whitened the data iteratively. Therefore, we considered the LISA displacement noise and TM acceleration noise, but no prior knowledge of the GW signals included in the simulated data. Further details on the fitting algorithm and the data preparation are given in [8].
In [7], the LISASim simulator [21] was used to generate the interferometer data. This simulator does not yet include GW signals that we want to study here. Therefore, we used the LISANode simulator [22, 23], which computes the response of given GW signals in dependence on the orbital position of the SC. In LISANode, the length differences are given as frequency fluctuations, unlike LISASim, where they are given as length variations. To use the same code as in [7], we converted the frequency data into length data. Even though LISANode simulates anti-alias filters, we applied an additional low-pass filter to enhance the robustness. In this step, we downsampled the data from 4 Hz to 2 Hz The same jitter levels and shapes have been implemented in both simulators. Moreover, only minor differences exist for the instrument noise contributors, such that we find an overall comparable level of instrument noise in the data from both simulators. The jitter and noise shapes used in our work are summarized in appendix A. We also find no significant differences in the accuracy of the TTL coupling coefficient estimation and the level of residual noise after TTL coupling subtraction. A more detailed comparison of the results using either of the two simulators is given in [8].
@@ IV TTL coupling estimation for changing arm lengths
In figure 2 we show the deviations of the estimated coupling coefficients from the true value. We find small differences between the estimates in the different cases. All estimation errors are below 0.06 mm/rad and therefore below the requirement of 0.1 mm/rad. Note that the error bars (1 $\sigma$ standard deviations) do not cover the difference between the single estimates and the true values in most cases. We attribute these deviations to the correlations of the coupling coefficients, which are not included in the standard deviations. The variances and covariances for the estimation method used are presented in [7], along with further discussion in [7, 8].
To demonstrate the performance of the TTL noise subtraction, we plot the residuals in figure 3. We find that the residuals after subtraction in the cases of the static (light-red) and the ESA science orbit (blue) are almost identical. We did not plot the residual for the Keplerian orbit case since it looks like the other two residual curves. This finding supports the assumption in [7] that it is sufficient to consider static unequal arm lengths in TTL coupling subtraction analyses. We also plot the TTL noise residual (green) for the scenario with ESA science orbits. In the present paper, the term TTL noise residual describes the non-suppressed TTL noise and the DWS noise added due to the subtraction [7, cf. eq. (18)]. It is well below the other instrument noises.
@@ V.1 Verification binaries
As an initial test of the TTL coupling coefficient estimation in the presence of GW signals, we include the response of two verification binaries. The signal characteristics of such binary systems are well studied and can be easily distinguished from the TTL coupling noise due to their distinct response at a single frequency. Therefore, we use them to verify our correct inclusion of the signals in the simulated data. We chose the verification binaries HMCnc and ZTFJ1539 for our analysis. Their gravitational wave frequencies are 6.220 mHz (HMCnc) and 4.822 mHz (ZTFJ1539) [31]. While most of the verification binaries listed in [31] enter the measurement band below the frequency range where TTL coupling is expected to dominate, the selected binaries lie in the frequency regime considered in our fit.
Figure 4: Performance of the TTL coupling subtraction in the scenario with verification binaries. The ASD of the TDI X combination for the simulations with TTL coupling is plotted in black. The dark-blue curve shows the ASD of the noise residual after the subtraction of the TTL coupling using the MCMC fit results and the light-blue dashed curve for the subtraction using the true coupling coefficients. The residual TTL coupling is plotted in dash-dotted green. The ASD of the GW response in TDI X is shown in purple. All ASDs were computed for one day measurements at 2 Hz. The dashed gray line corresponds to the LISA requirement.
@@ V.3 Detached white dwarf galactic binaries
Figure 9: Close-up view on the performance of the TTL coupling subtraction in the case GB1. The ASD of the TDI X combination for the simulations with (without) TTL coupling is plotted in black (gray). The dashed dark-blue curve shows the ASD of the residual after the subtraction of the TTL coupling using the MCMC fit results. The ASD of the GW response after TDI is shown in purple. The ASD of the residual after subtracting the TTL model with the fitted MCMC coefficients from data without instrument noise is plotted in dashed yellow. All ASDs were computed for one day measurements at 2 Hz.
@@ VI Summary
In the present paper, we have investigated the performance of the TTL coupling suppression via subtraction in post-processing for different GW signals in the simulated data. The data were generated using the simulator LISANode, and we chose the same noise realizations in all scenarios for a better comparison of the different cases. We then used an MCMC algorithm to fit the TTL coupling coefficients to the TDI AET combinations of the interferometric length measurements. Our analysis showed that the coupling coefficients could be fitted to within a deviation of $\pm$0.1 mm/rad in all the scenarios. By subtracting the TTL noise model from the data using the fitted coupling coefficients, we could suppress the total noise below the LISA mission noise requirement. The residual TTL noise was shown to be well below the other instrument noises. Due to the small deviations of the coupling coefficients, the differences of the residuals after subtraction using the fitted or the true coupling coefficients are marginal. We also investigated the GW signals after the TTL coupling subtraction. We could show that neither the coefficient inaccuracies nor the DWS readout noise added significantly to the data when subtracting the TTL noise. These results are a first evidence that the TTL suppression scheme planned for the LISA mission will not only suppress the noise well below the requirement, but will also not affect the GW signal content in the data. However, the latter was shown only indirectly in figures 9 and 14 for one day of data. The results would need to be confirmed considering the longer integration times for GW analyses and their ultimate effect on the astrophysical parameter estimation. Despite the overall reassuring results of our analysis, some aspects will require further investigation in the future. Our simulations of the LISA data considered only the SC and MOSA jitter, which is white in the frequency range where TTL coupling is assumed to be dominant. In reality, we expect the jitter to be colored. For instance, the SC jitter will roll off at higher frequencies reducing the frequency range in which TTL coupling is dominant. We have tested the TTL coefficient fit for smaller frequency ranges than the one used in most of this paper and found comparable results (see e.g. appendix B). It remains to be shown that the GW signals and the jitter can also be well distinguished for other jitter shapes. The implementation of the LISA dynamics and control loops [38, 39] in the LISANode simulator will allow more realistic tests in the future. Furthermore, our best estimates of instrument noise are constantly changing and may require updates of some of the presented tests. Finally, we have not yet considered all possible GW source types. E.g., we have not yet investigated extreme mass ratio inspirals or the cosmological SGWB.